OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66

OpenStudy (campbell_st):

well to complete this question you need to find the common difference before finding the 42nd term the formula for a term in an arithmetic sequence is \[T_{n} = a + (n - 1)d \] a = 1st term, n = number of terms and d is the common difference so using the known information you can say \[66 = -12 + (27 -1) d\] you need to solve for d.. \[66 = -12 + 26d\] when you have d, using the formula above with n = 42 to find the 42nd term hope it helps